Sir : An element with mass number 81 contains 31.7% more neutron than proton.Assign atomic Symbole ?

Grade:11

1 Answers

Vikas TU
6 years ago
MASS NO = 81 
p + n = 81 
in the event that prtons are x thn neutrons = ( x + 31.7/100) x = 1.317 x 
x + 131.7x = 81 , x = 35 , 
protons 35 implies nuclear number 35 , henceforth component is boron .